OpenStudy (somegirl1999):

Find the missing length. Round to the nearest tenth, if necessary. (1 point) 441 22.5 64 19.4

OpenStudy (somegirl1999):



OpenStudy (e.cociuba):

If c denotes the length of the hypotenuse and a and b denote the lengths of the other two sides, the Pythagorean theorem can be expressed as the Pythagorean equation is a2 + b2 = c2 Here a = 8, b = x and c = 21 So, 82 + x2 = 212 64 + x2 = 441 Subtract 64 from each side. x2 = 377 Take square root each side. x = 19.4 Therefore answer is x = 19.4

OpenStudy (somegirl1999):

thank you so much :)

OpenStudy (e.cociuba):

np:)